OpenStudy (anonymous):

convert the numeral to a numeral in base 10....1000_two.

OpenStudy (anonymous):

\[1000_{two} \]

OpenStudy (campbell_st):

so your number is \[1 \times 2^3 + 0 \times 2^2 + 0 \times 2^1 + 0 \times 2^0 = \] you just need to evaluate to get the answer.

OpenStudy (anonymous):

how do you even start to solve this lol

OpenStudy (anonymous):

or how did you get to what you got?

OpenStudy (anonymous):

if i evaluate that i get 8? is that right?

OpenStudy (campbell_st):

thats correct.

OpenStudy (anonymous):

what i dont get is how did you get what you got?

OpenStudy (campbell_st):

well its exactly the same as the decimal system